OpenStudy (anonymous):

3(6t-7)=2(9t-6) step by step plz

OpenStudy (anonymous):

3(6t-7)=2(9t-6) 18t-21=18t-12 0-21=-12 0=9 \[0\neq9\] DNE. No solutions exist for this.

OpenStudy (anonymous):

Step by step... \[3(6t-7)=2(9t-6)\]\[3\times 6t - 3\times 7 = 2\times 9t - 2 \times 6\]\[18t - 21 = 18t -12\]\[18t-21+21 = 18t - 12 + 21\]\[18t = 18t + 9\]\[0 \neq 9\]